OpenStudy (anonymous):

a car with a tire of radius 16 inches rotates at 450 revolution per minute. Find the speed of the car to the nearest kilometer per hour

OpenStudy (anonymous):

find the circumference multiply the circumference with the # revolutions per min and that will give you the speed in inches/min

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

2pi16=100.53

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

\[speed=\frac{ 450\times16\times2\pi \times 1.609344 km }{ \frac{ 1hour }{60 } }=\]

OpenStudy (anonymous):

why do i multiply 1.609344 km???

OpenStudy (anonymous):

well anyways i get 1213.42

OpenStudy (anonymous):

@phi

OpenStudy (phi):

this problem has two parts to it. First, you figure out that the circumference is 100.53 inches if it spins 450 times per minute, it will go 100.53 inches * 450 turns/1 minute = 45238.93 inches per minute that is a speed, but they want the units to be kilometers per hour we could use 1 inch = 2.54 cm write that as a ratio with inches on the bottom and multiply: \[ 45238.93 \cancel{\text{ in}} \cdot \frac{2.54 \text{ cm}}{1 \cancel{\text{ in}} }= 114,906.89 \text{ cm}\] now use 100 cm = 1 m multiply both sides by 1000 to get 100000 cm = 1000 m or 100000 cm = 1 km write this as a ratio with km on top and cm below, and multiply \[ 114,906.89 \cancel{\text{cm}} \cdot \frac{1 \text{ km}}{100000 \cancel{\text{cm}} } = 1.149 \text{ km}\] you now have 1.149 km per minute we now change minutes to hours using 60 min = 1 hr make a ratio with hr on top and min on the bottom \[ 1 \cancel{ \text{ min}} \cdot \frac{ 1 \text{ hr} }{60 \cancel{ \text{ min}}}= \frac{1 }{60} \text{ hr}\] you now have \[ \frac{ 1.149 \text{ km}}{\frac{1}{60} \text{ hr}} \] multiply top and bottom by 60 to get 68.94 km/hr round to the nearest kilometer per hour 69 km/hr